Levine Inc., which produces a single product, has prepared the following standard cost sheet for one unit of the product.

Direct materials (7 pounds at $1.60 per pound) $11.20
Direct labor (1 hours at $10.00 per hour) $10.00


During the month of April, the company manufactures 350 units and incurs the following actual costs.

Direct materials purchased and used (2,000 pounds) $3,600
Direct labor (370 hours) $3,663


Compute the total, price, and quantity variances for materials and labor.

Total materials variance $enter a dollar amount select a type of the variance FavorableUnfavorableNeither favorable nor unfavorable
Materials price variance $enter a dollar amount select a type of the variance FavorableNeither favorable nor unfavorableUnfavorable
Materials quantity variance $enter a dollar amount select a type of the variance UnfavorableNeither favorable nor unfavorableFavorable
Total labor variance $enter a dollar amount select a type of the variance FavorableNeither favorable nor unfavorableUnfavorable
Labor price variance $enter a dollar amount select a type of the variance FavorableUnfavorableNeither favorable nor unfavorable
Labor quantity variance $enter a dollar amount select a type of the variance FavorableNeither favorable nor unfavorableUnfavorable
Answer #1

Compute the total, price, and quantity variances for materials and labor.

Total materials variance (350*11.20-3600) = 320 Favorable
Materials price variance (1.6*2000-3600) = 400 Unfavorable
Materials quantity variance (350*7-2000)*1.6 = 720 Favorable
Total labor variance (3500-3663) = 163 Unfavorable
Labor price variance (10*370-3663) = 37 Favorable
Labor quantity variance (350-370)*10 = 200 Unfavorable